# “Set It Up!”: Functional Object Arrangement with Compositional Generative Models

### v1

@@ IV-B Evaluation Metrics
A orderly functional object arrangement must meet several criteria: physical feasibility (i.e., being collision-free), functionality (i.e., serving the intended purpose as per human instructions), and overall convenience and aesthetic appeal (i.e., neat arrangement with alignments and symmetries). While physical feasibility and functionality can be objectively evaluated using rule-based scripts, the subjective nature of convenience and aesthetic appeal precludes such fixed evaluations. Therefore, we employ both rule-based auto-grading and human experiments to evaluate models.
@@
5 Fully functional, and aesthetically pleasing. All items are well-placed, easy to access, and efficiently organized. Furthermore, the setup is visually appealing with proper alignment and symmetry.
@@ V-A Object Rearrangement
V-B Orderly Functional Object Arrangement
@@ A-B1 Task Prompt with Spatial Relationship Library
We use the same task prompt across all task families to guide the LLM in 1) proposing object relationships for orderly functional object arrangement (OFOA), and 2) presenting all available relationships from the relationship family $\mathcal{R}$. Crucially, each relationship is named to reflect its geometric property and includes a description to assist the LLM in understanding each relationship, facilitating the composition of the final constraint graph.
1 I would like to create an arrangement of objects that is both orderly (physically feasible and visually appealing) and functional (meets the users intended purpose as per instruction), referred to as OFOA, based on an under-specified user instruction. I need assistance in developing a Python rule template for generating these orderly functional object arrangements for <TASK FAMILY> using only the abstract object relationships
@@ 3 We provide the following library of abstract relationships:
53The orderly functional object arrangement for a <TASK FAMILY> should be grounded strictly using these abstract object relationships.
@@
55Your task is to create a rule template for OFOA (Orderly Functional Arrangement) applicable to <TASK FAMILY> in Python syntax. This template will summarize and abstract the common rules and patterns necessary for OFOA generation. You will be provided with 5 examples of orderly functional arrangements for <TASK FAMILY> alongside a sketch of a human-designed program. Please ensure you receive all examples and the program sketch before you start crafting the rule template.
@@ A-B2 Few-shot Examples
For each task family, we input five examples of the OFOA arrangement. Each example includes i) a user instruction outlining desired object arrangements, ii) the list of objects present on the table, and iii) the active abstract relationships. Objects are referred to by their types, allowing the LLM to deduce their functions. Abstract relationships are annotated by the abstract rule classifiers, as detailed in Appendix A-C, based on objects’ shapes and poses.
@@ 16 Returns:
Finally, we prompt LLM to generate the task-family OFOA rule template based on the program sketch and a few examples.
@@ A-B4 Self-reflective & CoT Output Prompt
4 2. Fully instantiate the rule template based on the given instruction and the object list, grounding the OFOA arrangement using the 24 defined abstract relationships.
@@ 20 Object list: [Insert object list here]
22 Your task is to arrange these objects using the 24 types of abstract relationships to create an OFOA (Orderly Functional Arrangement). Extract all proposed relationships into a list, following this format:
@@ A-E Baselines
End-to-end Diffusion Model: Drawing inspiration from StructDiffusion [18], we developed a language-conditioned diffusion model. This model is designed to predict object poses in OFOA arrangements. It has been trained using the same datasets as our main model. These datasets include a synthetic object relationship dataset and 15 specific OFOA examples.
For synthetic objects, we employ a script that describes the geometric relationships between scene objects as the language input. For OFOA scenes, the language input comes from user instructions. We then encode these language inputs into language feature embedding using CLIP embedding. In our model, each object is represented by a combined token of its shape and pose embeddings. We have configured the transformer to handle up to 16 objects, which corresponds to its sequence length limit of 16. The model processes the language embeddings together with object sequence embeddings and time embeddings through a transformer encoder. This encoder uses four layers of residual attention blocks and is equipped with two heads. We also incorporate normalization layers both before and after the transformer layers. Finally, we utilize the same diffusion timesteps $t=1500$ in this model as in our SetItUp.
@@
Direct LLM prediction Drawing on Tidybot’s approach, we use an LLM-based method to directly propose the object poses for OFOA arrangements. Specifically, we feed the LLM the same task instruction, along with a library of relationships and the few-shot examples. Additionally, we provide details of the table layout and include the object shapes in the few-shot training examples. It’s important to note that, unlike our model, which only processes abstract relationships, the LLM is given actual shapes from the ground truth. We then instruct the LLM to identify common patterns within these examples, following the methodology proposed by Tidybot. For inference, we supply the user instructions, a list of objects, shapes of these objects to the LLM. Then, we ask the LLM to predict the poses for objects in OFOA arrangements.
@@ 12 To propose the tidy arrangements, strictly follow the following structure:
  • Pose Inference: Tidybot requires the LLM to output the abstract action primitives and the receptables for “tidy arrangement”, yet we need the object poses. Therefore, while the tasks in the Tidybot project requires no geometric understanding from LLM, OFOA requires the LLM to be able to understand the geometries so as to propose the object poses. To help LLM with this, we provide the table layout and the object dimensions for few-shot learning.
@@ 7 Object list: [Insert object list here]
9 Your task is to arrange these objects using the 24 types of abstract relationships to create an OFOA (Orderly Functional Arrangement). Extract all proposed relationships into a list, following this format:
@@ A-F1 Study Desk
  • Training & Testing datasets: Table VI contains the training examples, each consisting of user instructions, object lists, and OFOA layouts. The testing dataset, which includes user instructions and object lists, is presented in Table VII.
@@
Index User Instruction Object List OFOA Arrangement
@@
Index User Instruction Object List OFOA Arrangement
@@ A-F2 Coffee Table
Index User Instruction Object List OFOA Arrangement
@@
  • Training & Testing datasets: Table VIII contains the training examples, each consisting of user instructions, object lists, and OFOA layouts. The testing dataset, which includes user instructions and object lists, is presented in Table IX.
@@ A-F3 Dining Table
  • Training & Testing datasets: Table X contains the training examples, each consisting of user instructions, object lists, and OFOA layouts. The testing dataset, which includes user instructions and object lists, is presented in Table XI.
@@ 15 Returns:
Step 1: Task-Relevant Information: Users determine which information significantly impacts object-functional arrangement (OFOA). This involves writing function specifications to query this information.
@@
Step 2: Functional Clustering of Objects: Define key categories for clustering objects for OFOA. Write function specifications to classify input objects into these categories.

### v3

@@ IV-B Evaluation Metrics
A functional object arrangement must meet several criteria: physical feasibility (i.e., being collision-free), functionality (i.e., serving the intended purpose as per human instructions), and overall convenience and aesthetic appeal (i.e., neat arrangement with alignments and symmetries). While physical feasibility and functionality can be objectively evaluated using rule-based scripts, the subjective nature of convenience and aesthetic appeal precludes such fixed evaluations. Therefore, we employ both rule-based auto-grading and human experiments to evaluate models.
@@
5 Fully functional, and aesthetically pleasing. All items are well-placed, easy to access, and efficiently organized. Furthermore, the setup is visually appealing with proper alignment and symmetry.gra
@@ V-A Object Rearrangement
V-B Functional Object Arrangement
@@ VI Conclusion
Acknowledgments
This research is supported in part by the National Research Foundation (NRF), Singapore and DSO National Laboratories under the AI Singapore Program (AISG Award No: AISG2-RP-2020-016), NSF grant 2214177, AFOSR grant FA9550-22-1-0249, ONR MURI grant N00014-22-1-2740, and ARO grant W911NF-23-1-0034. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of NRF Singapore.
@@ A-B1 Task Prompt with Spatial Relationship Library
We use the same task prompt across all task families to guide the LLM in 1) proposing object relationships for functional object arrangement (FORM), and 2) presenting all available relationships from the relationship family $\mathcal{R}$. Crucially, each relationship is named to reflect its geometric property and includes a description to assist the LLM in understanding each relationship, facilitating the composition of the final constraint graph.
1 I would like to create an arrangement of objects that is both orderly (physically feasible and visually appealing) and functional (meets the users intended purpose as per instruction), referred to as FORM, based on an under-specified user instruction. I need assistance in developing a Python rule template for generating these functional object arrangements for <TASK FAMILY> using only the abstract object relationships
@@ 3 We provide the following library of abstract relationships:
53The functional object arrangement for a <TASK FAMILY> should be grounded strictly using these abstract object relationships.
@@
55Your task is to create a rule template for FORM (Functional Object Arrangement) applicable to <TASK FAMILY> in Python syntax. This template will summarize and abstract the common rules and patterns necessary for FORM generation. You will be provided with 5 examples of functional arrangements for <TASK FAMILY> alongside a sketch of a human-designed program. Please ensure you receive all examples and the program sketch before you start crafting the rule template.
@@ A-B2 Few-shot Examples
For each task family, we input five examples of the FORM arrangement. Each example includes i) a user instruction outlining desired object arrangements, ii) the list of objects present on the table, and iii) the active abstract relationships. Objects are referred to by their types, allowing the LLM to deduce their functions. Abstract relationships are annotated by the abstract rule classifiers, as detailed in Appendix A-C, based on objects’ shapes and poses.
@@ 16 Returns:
Finally, we prompt LLM to generate the task-family FORM rule template based on the program sketch and a few examples.
@@ A-B4 Self-reflective & CoT Output Prompt
4 2. Fully instantiate the rule template based on the given instruction and the object list, grounding the FORM arrangement using the 24 defined abstract relationships.
@@ 20 Object list: [Insert object list here]
22 Your task is to arrange these objects using the 24 types of abstract relationships to create an FORM (Orderly Functional Arrangement). Extract all proposed relationships into a list, following this format:
@@ A-E Baselines
End-to-end Diffusion Model: Drawing inspiration from StructDiffusion [18], we developed a language-conditioned diffusion model. This model is designed to predict object poses in FORM arrangements. It has been trained using the same datasets as our main model. These datasets include a synthetic object relationship dataset and 15 specific FORM examples.
For synthetic objects, we employ a script that describes the geometric relationships between scene objects as the language input. For FORM scenes, the language input comes from user instructions. We then encode these language inputs into language feature embedding using CLIP embedding. In our model, each object is represented by a combined token of its shape and pose embeddings. We have configured the transformer to handle up to 16 objects, which corresponds to its sequence length limit of 16. The model processes the language embeddings together with object sequence embeddings and time embeddings through a transformer encoder. This encoder uses four layers of residual attention blocks and is equipped with two heads. We also incorporate normalization layers both before and after the transformer layers. Finally, we utilize the same diffusion timesteps $t=1500$ in this model as in our SetItUp.
@@
Direct LLM prediction Drawing on Tidybot’s approach, we use an LLM-based method to directly propose the object poses for FORM arrangements. Specifically, we feed the LLM the same task instruction, along with a library of relationships and the few-shot examples. Additionally, we provide details of the table layout and include the object shapes in the few-shot training examples. It’s important to note that, unlike our model, which only processes abstract relationships, the LLM is given actual shapes from the ground truth. We then instruct the LLM to identify common patterns within these examples, following the methodology proposed by Tidybot. For inference, we supply the user instructions, a list of objects, shapes of these objects to the LLM. Then, we ask the LLM to predict the poses for objects in FORM arrangements.
@@ 12 To propose the tidy arrangements, strictly follow the following structure:
  • Pose Inference: Tidybot requires the LLM to output the abstract action primitives and the receptables for “tidy arrangement”, yet we need the object poses. Therefore, while the tasks in the Tidybot project requires no geometric understanding from LLM, FORM requires the LLM to be able to understand the geometries so as to propose the object poses. To help LLM with this, we provide the table layout and the object dimensions for few-shot learning.
@@ 7 Object list: [Insert object list here]
9 Your task is to arrange these objects using the 24 types of abstract relationships to create an FORM ( Functional Object Arrangement). Extract all proposed relationships into a list, following this format:
@@ A-F1 Study Desk
  • Training & Testing datasets: Table VI contains the training examples, each consisting of user instructions, object lists, and FORM layouts. The testing dataset, which includes user instructions and object lists, is presented in Table VII.
@@
Index User Instruction Object List FORM Arrangement
@@
Index User Instruction Object List FORM Arrangement
@@ A-F2 Coffee Table
Index User Instruction Object List FORM Arrangement
@@
  • Training & Testing datasets: Table VIII contains the training examples, each consisting of user instructions, object lists, and FORM layouts. The testing dataset, which includes user instructions and object lists, is presented in Table IX.
@@ A-F3 Dining Table
  • Training & Testing datasets: Table X contains the training examples, each consisting of user instructions, object lists, and FORM layouts. The testing dataset, which includes user instructions and object lists, is presented in Table XI.
@@ 15 Returns:
Step 1: Task-Relevant Information: Users determine which information significantly impacts functional object arrangement (FORM). This involves writing function specifications to query this information.
@@
Step 2: Functional Clustering of Objects: Define key categories for clustering objects for FORM. Write function specifications to classify input objects into these categories.